### v1

# Dialogue Without Limits: Constant-Sized KV Caches for Extended Responses in LLMs

Abstract
Autoregressive Transformers rely on Key-Value (KV) caching to accelerate inference. However, the linear growth of the KV cache with context length leads to excessive memory consumption and bandwidth constraints. This bottleneck is particularly problematic in real-time applications – such as chatbots and interactive assistants – where low latency and high memory efficiency are critical. Existing methods drop distant tokens or compress states in a lossy manner, sacrificing accuracy by discarding vital context or introducing bias.
We propose MorphKV, an inference-time technique that maintains a constant-sized KV cache while preserving accuracy. MorphKV balances long-range dependencies and local coherence during text generation. It eliminates early-token bias while retaining high-fidelity context by adaptively ranking tokens through correlation-aware selection. Unlike heuristic retention or lossy compression, MorphKV iteratively refines the KV cache via lightweight updates guided by attention patterns of recent tokens. This approach captures inter-token correlation with greater accuracy, crucial for tasks like content creation and code generation. Our studies on long-response tasks show 52.9% memory savings and 18.2% higher accuracy on average compared to state-of-the-art prior works, enabling efficient real-world deployment.
Machine Learning, ICML
1 Introduction
Large Language Models (LLMs) have become indispensable for tasks requiring extensive context retention (e.g., document summarization) and prolonged text generation (e.g., code synthesis). As model architectures become sophisticated, their ability to process nuanced inputs and produce coherent, long-form outputs has improved dramatically. However, this progress is hindered by the memory overhead of Key-Value (KV) caches. KV caches store the key-value pairs to enable attention mechanisms for auto-regressive decoding during LLM inference. Unfortunately, as shown in Figure 1, the KV cache size grows with sequence length, often exceeding the memory capacity of even high-end GPUs.
The distinction between long-context and long-response tasks lies in the phase where token processing dominates. Long-context tasks, such as document summarization and prompt comprehension, primarily process a large volume of input tokens during the prefill phase, where the model ingests and encodes the initial prompt. In contrast, long-response tasks, such as essay writing and code generation, generate a substantial number of output tokens during the decode phase, requiring sustained attention over growing sequences of self-generated tokens.
Figure 1: KV cache sizes for the Llama 3.1 70B and 405B models across varying sequence lengths with a batch size of 256.
Figure 2: (a) Despite compression, the state-of-the-art SnapKV memory footprint increases with response length and exceeds available HBM capacity even on high-end GPUs. This study uses the Qwen 2.5 7B model on an NVIDIA H100 and the LongWriter benchmark. (b) Even at lower memory capacity, MorphKV achieves higher accuracy than SnapKV for long-response tasks.
Figure 3: Illustrative comparison of KV cache reduction methods as tokens are processed. (a) Scissorhands retains only a window of recent (shown in green) tokens, (b) StreamingLLM also stores a few initial tokens (old shown in yellow) from the prefill step, and (c) H2O stores even more old tokens (all prompt tokens) and only relevant recent tokens. (d) Keyformer stores only important old and recent tokens but remains biased towards the early tokens. (e) SnapKV retains selected old tokens from prefill and all decode (more recent) tokens (f) MorphKV identifies and stores only those old tokens that correlate with recent tokens.
Numerous approaches have been proposed in the literature to minimize the impact of growing KV cache sizes. In studies like FlashAttention (Dao, 2023) and vLLM (Kwon et al., 2023), the authors propose techniques to either materialize only partial caches at a time or use paging techniques by fragmenting KV caches into smaller blocks, thereby avoiding the need to reserve memory for the entire cache at once. Beyond this, prior works like FastGen (Ge et al., 2023), and MiKV (Yang et al., 2024) compress KV caches by retaining only a subset of KV pairs from recent and older tokens, prioritizing those deemed important based on attention scores and discarding the rest. However, this creates a trade-off: while memory savings increase as more KVs are discarded, the accuracy depends on the retained KVs effectively capturing context for future tokens. Consequently, these methods often sacrifice accuracy for reduced memory usage.
For example, as shown in Figure 3, Scissorhands (Liu et al., 2024b) retains only the KVs of recent tokens, sacrificing accuracy by discarding past context. StreamingLLM (Xiao et al., 2023) improves accuracy slightly by preserving KVs of a few initial tokens (attention sinks) alongside recent tokens but struggles when early tokens fail to capture sufficient context. H2(Zhang et al., 2023) retains KVs from the entire input prompt and the most attended output tokens, achieving high performance but reduced memory savings due to the large number of retained KVs. It also suffers from selection bias during decoding, preserving unimportant past KVs, which hinders performance in long-response tasks.
Keyformer (Adnan et al., 2024) selects top old and recent tokens for memory savings and uses Gumbel noise to reduce selection bias. While somewhat effective, it cannot fully eliminate selection bias because Gumbel noise by itself introduces new forms of biases (Mussmann et al., 2017), reducing accuracy for long-context and long-response tasks. SnapKV (Li et al., 2024), the current state-of-the-art, achieves high accuracy in long-context tasks by retaining the most attended tokens from the input prompt. However, as shown in Figure 2(a), it retains all generated output tokens from the decode phase, causing the KV cache size to scale with response lengths, making it unsuitable for long-response tasks. Overcoming these limitations is crucial for improving LLM performance in applications like scripting and content creation (long-response tasks).
Our Proposal – MorphKV: MorphKV achieves a constant-size KV cache by retaining only a limited number of old and recent tokens. However, to achieve higher accuracy, MorphKV employs a more dynamic KV selection algorithm that analyzes the attention patterns of the current token toward retained KVs. Unlike prior methods that independently identify important tokens, MorphKV retains only those old tokens that correlate strongly with recent tokens.
To better capture context, MorphKV prioritizes the attention scores of relevant recent tokens rather than relying on historically most-attended tokens, addressing bias issues observed in methods like H2O and Keyformer. As shown in Figure 2(b), MorphKV achieves better scores than SnapKV: for Phi4, up to 8% higher for long-context task (VCSum) while saving 56% on KV cache memory, and for Qwen2.5, up to 21% higher score for long-response task (LongGenBench) while saving 83% on KV cache memory. This shows the impact of retaining a compact set of high-quality KVs and an improved attention mechanism proposed in MorphKV.
MorphKV improves accuracy by 9.4% and 18.2% on average compared to SnapKV and H2O while reducing the KV cache footprint by 88.1% and 52.9% respectively for long-response tasks.
2 Background and Motivation
2.1 Large Language Model Inference
LLM inference begins with the prefill step, where the model processes the input prompt and generates Key-Value (KV) pairs for each token in the prompt. Next, in the decode phase, the model generates output tokens auto-regressively such that each output token attends to the KV pairs of all preceding tokens in the sequence while creating its own KV pair. This attention mechanism enables the LLM to maintain context and produce coherent responses. The KV pairs are stored in memory structures known as KV caches.
KV caches scale with the number of tokens processed, becoming prohibitively large for long-context and long-response tasks and posing significant challenges in deploying LLMs. Long-context tasks, such as creating diet plans from medical histories or summarizing documents like manuals, loan agreements, or papers, involve long prompts with many input tokens. In contrast, long-response tasks such as crafting lesson plans, providing step-by-step instructions, or writing scripts generate numerous output tokens from short inputs. While both types of tasks require large KV caches, they differ in when the KVs are produced. Long-context tasks generate most KVs in the prefill step, unlike long-response tasks that create most KVs during decoding.
2.2 Limitations of KV Cache Compression Methods
KV cache compression addresses their growing memory footprint through quantization, algorithmic optimizations, cross-layer approaches, and pruning. Quantization-based methods store KV pairs using lower precision (Kang et al., 2024; Zhang et al., 2024), whereas algorithmic methods modify attention-layer computations (Chang et al., 2024; Saxena et al., 2024). Cross-layer optimizations leverage inter-layer similarities, selectively retaining KVs from layers with significant contributions (Yuan et al., 2024; Saxena et al., 2024; Cai et al., 2024). Pruning strategies selectively retain a subset of important KV pairs (Adnan et al., 2024; Xiao et al., 2023; Li et al., 2024; Zhang et al., 2023; Liu et al., 2024b) and are more effective than other approaches because they capture task-specific context more accurately.
However, compressed KV caches are either limited by accuracy or scalability. Constant-sized KV caching methods, such as Scissorhands, StreamingLLM, and Keyformer, have limited accuracy. On the other hand, more accurate methods like SnapKV are not scalable as they fail to address growing KV cache size for long-response tasks.
3 MorphKV
This paper proposes MorphKV, a KV compression technique that reduces the KV cache size without compromising accuracy. MorphKV partitions the context into two components: recent context ($\mathcal{R}$) and distant context ($\mathcal{D}$). The recent context $\mathcal{R}$ corresponds to the last $R$ tokens that preserve local coherence, while the distant context $\mathcal{D}$ captures long-range dependencies. By attending to both $\mathcal{R}$ and a selective subset of $\mathcal{D}$, MorphKV ensures that the generated text remains contextually coherent and semantically meaningful.
Figure 4 presents an overview of our proposed design. A key insight in MorphKV is that tokens in $\mathcal{R}$ have already attended to tokens in $\mathcal{D}$ during their generation. Therefore, rather than retaining all or a subset of older tokens based on aggregated patterns, MorphKV leverages the attention profiles of recent tokens to select only the most relevant distant tokens. In this way, MorphKV constructs a compact yet accurate KV cache of size $C+R$, where $C$ is the number of distant tokens retained and $R$ is the number of recent tokens. Specifically, MorphKV 1 ranks older tokens based on their relevance to the recent tokens using a specialized algorithm that performs element-wise transformations, denoted by $f(x)$ in Figure 4. As attention scores inherently quantify how strongly past tokens were attended to during prior generations, using the attention scores of the recent tokens helps surface the most contextually relevant older tokens. Next, MorphKV 2 selectively retains only the most correlated old tokens in the KV cache, evicting those deemed irrelevant. This approach ensures an optimized memory footprint while preserving essential long-range dependencies.
Figure 4: Overview of MorphKV. (1) MorphKV uses the most recent window tokens to capture neighboring context and their attention scores to rank the older tokens. (2) To capture relevant distant context, MorphKV only retains old tokens maximally correlated to the recent window tokens by consulting the attention scores aggregated using a fusion function $f(x)$.
3.1 Mathematical Formulation
Let $Q_{i}$, $K_{i}$, and $V_{i}$ be the query, key, and value vectors for the token being generated at timestep $i$. Let $G_{i}$ denote the KV cache storing $(K_{j},V_{j})$ pairs for all the previously generated tokens $j<i$. The standard attention mechanism (Vaswani et al., 2023) computes the attention weights $AW_{i}$, as shown in Equation (1):
$AW_{i}\;=\;\text{Softmax}\!\Bigl{(}\frac{Q_{i}\,K^{T}}{\sqrt{d_{h}}}\Bigr{)},\quad O_{i}\;=\;AW_{i}\cdot V,$ (1)
where $K=[K_{1},K_{2},\dots,K_{i-1}]\in\mathbb{R}^{(i-1)\times d_{h}}$, $V=[V_{1},V_{2},\dots,V_{i-1}]\in\mathbb{R}^{(i-1)\times d_{h}}$, and $d_{h}$ is the hidden dimension. The attention output $O_{i}\in\mathbb{R}^{d_{h}}$ encodes the information from all the previous tokens.
3.1.1 Problem Statement
Retaining every key-value pair $(K_{j},V_{j})$ for $j<i$ can increase memory usage as $i$ grows. Let $G_{i}^{*}$ be an optimal reduced cache of size $C+R$ that minimizes the change in the attention output as denoted in Equation (2):
$G_{i}^{*}\;=\;\arg\min_{\begin{subarray}{c}G_{i}^{\prime}\subseteq G_{i}\\ |G_{i}^{\prime}|=C+R\end{subarray}}\bigl{\|}\,O_{i}-O_{i}^{\prime}\bigr{\|}_{2},$ (2)
where $O_{i}^{\prime}$ is the attention output, as shown in Equation 1, computed using only the tokens in $G_{i}^{\prime}$. We want $\|O_{i}-O_{i}^{\prime}\|_{2}\leq\epsilon$ for a small $\epsilon\geq 0$, ensuring minimal error despite reducing the KV cache to $C+R$ entries.
3.1.2 Approximating Optimal KV cache
Figure 5: Illustration of the Key Value (KV) caching mechanism in MorphKV. MorphKV uses the insight that recent tokens naturally capture some distant context from old tokens due to the auto-regressive nature of token generation. For example, at decoding step $T_{0}$, MorphKV consults the Attention Profile of recent tokens weather and The to learn that these tokens have attended considerably more to the old token today’s than me. MorphKV uses this information and evicts the latter.
Although solving Equation 2 directly is intractable because of its combinatorial nature, MorphKV adopts two intuitive heuristics (H1 and H2):
  • Local Coherence (H1): Always retain the last $R$ tokens to preserve continuity.
  • Distant Relevance (H2): Retain only the $C$ most informative older tokens, as measured by fused attention scores with respect to the $R$ recent tokens.
Concretely, we define an approximate policy $\mathcal{P}^{\prime}$ in Equation (3) as shown below:
$G_{i}^{\prime}\;=\;\mathcal{P}^{\prime}\bigl{(}G_{i},F_{i}\bigr{)},\quad\text{where}\;\;|G_{i}^{\prime}|=C+R.$ (3)
Here, $G_{i}^{\prime}\subseteq G_{i}$ contains (1) the $R$ most recent entries $G_{i}^{R}$ and (2) the top-$C$ older entries selected based on an auxiliary score vector $F_{i}$.
3.1.3 Developing the Auxiliary Score Vector $F_{i}$
Let $\mathcal{W}_{i}=\{w_{i-1},w_{i-2},\dots,w_{i-R}\}$ denote the window of $R$ most recent tokens at timestep $i$. MorphKV inspects the attention weights of these tokens to build $F_{i}$, using Equation (4). Specifically,
$F_{i}[k]\;=\;f\bigl{(}AW_{i-1}[k],\,AW_{i-2}[k],\,..,\,AW_{i-R}[k]\bigr{)}$ (4)
where $AW_{r}[k]$ is the attention weight that $w_{r}$ assigned to the $k$-th older token, and $f(\cdot)$ is a fusion function. For $f(\cdot)$, MorphKV proposes two choices, the sum and max fusions, as shown in Equations (5) and (6) respectively:
Sum Fusion: $F_{i}[k]\;=\;\sum_{r=i-1}^{\,i-R}\!AW_{r}[k]$ (5)
Max Fusion: $F_{i}[k]\;=\;\max_{r=i-1}^{\,i-R}\!AW_{r}[k].$ (6)
The Sum Fusion prefers tokens consistently attended to across multiple recent tokens, whereas the Max Fusion selects tokens strongly attended by at least one recent token. The intuition is that tokens frequently or strongly attended to by recent tokens are likely critical for maintaining long-range coherence. For example, recurring entities, such as characters in a story, often receive sustained attention across multiple decoding steps. Algorithm 1 shows the dynamic token selection process using the auxiliary score vector.
Algorithm 1 MorphKV: Dynamic Token Selection
  Input:
  • KV cache $G_{i}$ (keys/values of older tokens)
  • Tokens $x_{1:i}$ (sequence generated so far)
  • Num of Query-Heads, Key-Heads: $\{M,M^{\prime}\}$
  • Per-head attention weights $\{AW^{m}\}_{m=1}^{M}$
  • Window size $R$, fusion function $f$, capacity $C$
  Output: Updated cache $G_{i+1}$
  $\mathcal{W}_{i}\leftarrow\{x_{i-1},x_{i-2},\dots,x_{i-R}\}$triangleright Recent tokens
  for $w_{r}\in\mathcal{W}_{i}$ do
     $S_{r}\leftarrow\sum_{m=1}^{(M/M^{\prime})}AW_{r}^{m}$triangleright Aggregate scores if GQA
  end for
  $F_{i}\leftarrow f(S_{i-1},S_{i-2},\dots,S_{i-R})$triangleright Fuse recent tokens’ scores
  $G_{i+1}\leftarrow\text{Top}_{C}(F_{i})\ \cup\ \mathcal{W}_{i}$triangleright Retain top-$C$ distant + $R$ recent
  Return: $G_{i+1}$
3.1.4 Selection of KV Pairs
After computing $F_{i}$, we pick the top-$C$ entries (older tokens) according to $F_{i}$ and combine them with the $R$ most recent tokens, as shown in Equation (7):
$G_{i+1}=\Bigl{\{}\!\text{Top}_{C}(F_{i})\Bigr{\}}\;\cup\;\Bigl{\{}\!\text{$R$ recent tokens}\Bigr{\}}$ (7)
Hence, $G_{i+1}$ contains $C+R$ tokens in total, satisfying heuristics (H1) and (H2). By updating $G_{i}\to G_{i+1}$ at each timestep, MorphKV prunes the KV cache incrementally, ensuring that memory usage remains fixed at $C+R$ while preserving essential local and distant context.
3.2 Intuition with A Walk-Through Example
Figure 5 demonstrates the operations in MorphKV with $C=2$ and $R=2$, using sum fusion as the fusion function.
At timestep $T_{0}$: Suppose the recent tokens are [weather, The], while the old tokens are [me, today’s]. The attention profiles (AP) of weather and The both strongly point to today’s (e.g., $0.3$ each) and weakly to me (e.g., $0.05$ each). By summing these attention scores, MorphKV deems today’s to be more relevant (score $0.6$) than me (score $0.1$). Consequently, it retains today’s in $\mathcal{D}$ and evicts me.
At timestep $T_{1}$: The most recent tokens shift to [perfect, for], while older tokens include [sun, out]. MorphKV recalculates the fused attention scores for sun and out. If out receives a lower combined score than sun (as illustrated), it is evicted, preserving only sun in the distant context.
At timestep $T_{2}$: The most recent tokens are now [for, bright], and the older tokens still include sun. If sun continues to receive relatively high attention from the new recent tokens, it remains in the cache despite being one of the oldest tokens. This shows how MorphKV can preserve distant tokens of continued importance (e.g., sun) while evicting those that have very likely become less relevant over time.
3.3 Handling Multiple Heads
For transformer-based LLMs, attention is performed across multiple heads, also referred to as Multi-Headed Attention (MHA)  (Vaswani et al., 2023). In MHA with $M$ heads, each head maintains a separate KV cache, resulting in $M$ distinct caches and significant memory overhead. Modern LLMs like Llama3.1  (Dubey et al., 2024) use Grouped-Query Attention (GQA) (Ainslie et al., 2023) with $M^{\prime}$ grouped-key heads ($M^{\prime}<M$), where every $M/M^{\prime}$ query heads share a single key/value head, reducing the number of KV caches to $M^{\prime}$ and cutting memory usage by a factor of $M/M^{\prime}$.
MorphKV is compatible with both MHA and GQA architectures. For MHA, we independently apply MorphKV to each of the $M$ heads, pruning their caches in parallel. For GQA, we first aggregate attention weights from the $M/M^{\prime}$ grouped query heads, then compute fusion scores $F_{i}$ for the shared key-value heads. Our experiments show that MorphKV performs equally well with both approaches. By default, we choose GQA due to its memory efficiency. This flexibility distinguishes MorphKV from prior works like SnapKV (Li et al., 2024) which are only limited to MHA.
4 Evaluation Methodology
Models: We evaluate MorphKV across four state-of-the-art LLMs chosen based on their complementary strengths:
  • Llama-3.1 8B Instruct (Dubey et al., 2024): A model optimized for long-context tasks (128K token window) and coherent multi-turn dialogue.
  • Mistral-v0.2 7B Instruct (Jiang et al., 2023): A lightweight architecture designed for efficient deployment on consumer hardware.
  • Qwen2.5 7B Instruct (Qwen et al., 2025): A model with multi-lingual English and Chinese proficiency.
  • Phi-4 14B (Abdin et al., 2024): A model specialized in STEM reasoning through high-quality training data and curriculum learning.
This diversity ensures rigorous validation of MorphKV’s robustness, scalability, and cross-architectural consistency.
Setup: We run experiments on an NVIDIA Grace Hopper node with an H200 GPU (96GB HBM3) and Grace CPU (116GB LPDDR5) interconnected via NVLink. We implement MorphKV using HuggingFace Transformers (Wolf, 2020) with FlashAttention-2 (Dao, 2023) for hardware-aware optimization, mirroring the configuration of prior KV cache works (Adnan et al., 2024; Li et al., 2024; Wang et al., 2024; Zhang et al., 2023).
Benchmarks:
  • Long-Response generation: LongWriter (Bai et al., 2024) and LongGenBench (Liu et al., 2024a), which require synthesizing structured outputs (e.g., diaries, floorplans) based on input prompts.
  • Long-Context understanding: LongBench (Bai et al., 2023), for tasks like code repository navigation and document summary with 16K-128K token contexts.
Baseline: We compare the performance and memory efficiency of MorphKV against SnapKV (Li et al., 2024), H2(Zhang et al., 2023), and Full-Attention. SnapKV is the state-of-the-art for KV cache compression, while Full-Attention provides an upper bound on accuracy. However, SnapKV does not perform token eviction during the generation phase, making it less efficient for long-response tasks where cache management is critical. H2O applies KV cache pruning and is thus a more meaningful baseline for evaluating MorphKV in long-response settings. Prior works retain KV pairs across all attention heads, while MorphKV’s compatibility with Grouped-Query Attention (GQA) enables a more memory-efficient approach as it only retains KV pairs across grouped-key heads, allowing us to assess trade-offs between KV cache size and retention of relevant tokens across different models and benchmarks.
Implementation: We implement MorphKV in HuggingFace transformers library (Wolf, 2020), integrating it into the existing attention mechanism and leveraging FlashAttention (Dao, 2023) for efficient inference. To extract attention scores for older tokens, we compute partial attention weights for window queries within FlashAttention and store them as a lightweight KV cache extension. We update this cache during generation by appending new attention profiles and discarding the oldest ones.
5 Results
5.1 Long-Response: LongWriter Tasks
We evaluate MorphKV on open-ended, long-response text generation using the LongWriter (en) benchmark. LongWriter covers tasks such as writing emails, blog posts, essays, and novels, with 60 prompts requesting responses ranging from 100 to 12000 words. For comparison, SnapKV retains 600 prompt tokens plus all decoded tokens across all attention heads, while H2O stores 600 decoded tokens per head. In contrast, MorphKV maintains a recent window of 30 tokens, a total KV cache capacity of 600 tokens, and supports two fusion strategies: sum() and max().
5.1.1 Performance
Performance is measured using an LLM-based Judge (Mistral-Large-123B), which assigns scores based on several criteria, and these scores are aggregated to form a final overall score. Table 1 shows that MorphKV outperforms both H2O, and SnapKV on Llama, Mistral and Phi4 models, while achieving comparable performance to SnapKV for Qwen. Notably, MorphKV consistently scores higher than H2O on metrics such as relevance (see Appendix A.2), indicating its ability to track important context for long responses when constrained by memory size.
xModelx Llama Mistral Phi4 Qwen
H2O 68.5 80.0 61.5 63.8
SnapKV 67.7 81.1 63.8 68.4
MorphKV 69.5 81.1 64.7 64.9
Full-Attention 66.5 81.3 62.9 66.2
Table 1: LongWriter: LLM Judge Score comparison of MorphKV, H2O, SnapKV, and Full-Attention scores across different models. MorphKV outperforms other techniques by upto $4.5\%$, retaining important older tokens at a much smaller memory footprint.
5.1.2 KV cache Sizes
Figure 6 shows the normalized KV cache sizes for H2O, SnapKV, and MorphKV relative to Full-Attention. On average, MorphKV needs only 0.25×0.25$\times$0.25 × the memory capacity of Full-Attention, while H2O needs roughly 1$\times$ and SnapKV needs as much as 4$\times$ the memory capacity of Full-Attention.
Figure 6: LongWriter: KV cache sizes for H2O, SnapKV, MorphKV relative to Full-Attention. On average,  MorphKV requires 0.25$\times$ the KV cache size used by Full-Attention, while H2O and SnapKV use 1$\times$ and 4$\times$ respectively.
5.1.3 Impact of Increasing Response Length
To evaluate the robustness of MorphKV for long responses, we compute the LLM Judge Score against increasing response lengths, as shown in Figure 7 for Mistral-7B. As the response length increases, performance declines across all methods due to the inherent challenges of generating extremely long text (Bai et al., 2024). However, MorphKV degrades more gradually: a 4×4$\times$4 × increase in length reduces performance by $15\%$$18\%$ for SnapKV and H2O, whereas the performance only reduces by $10\%$ for MorphKV. Notably, MorphKV maintains a constant KV cache size regardless of the response length, contributing to its efficiency and robustness over extended text generations.
Figure 7: LongWriter: LLM Judge Score versus response lengths. MorphKV is more robust against increasing response lengths compared to SnapKV or H2O as it uses an adaptive KV selection algorithm, discarding those KVs which don’t contribute significantly to the contextual flow. Notably, MorphKV maintains a fixed KV cache size regardless of response length.
5.2 Long-Response: LongGenBench Tasks
LongGenBench contains structured long-response tasks for temporal and spatial categories. The temporal category is divided into Diary Entry and Menu planning tasks, while the spatial category includes Skyscraper Design and Urban Planning. The original dataset has 400 samples (100 from each sub-category). For a fair comparison under limited resources, we select 40 samples, with ten from each sub-category, and use greedy decoding capped at 8K tokens. SnapKV employs all attention heads with a 32-token window and a total KV cache capacity of 1K tokens per head, whereas H2O keeps 4K tokens in the cache, also maintaining all attention heads. In contrast, MorphKV uses a 200-token recent window and a 4K-token total capacity across GQA heads, adopting max() fusion due to its consistently higher performance than sum(). The larger window enables MorphKV to retain critical distant tokens during generation (refer to Section 6.2 for more information).
Table 2: LongGenBench: Performance comparison of MorphKV, SnapKV, H2O across different models. MorphKV achieves better scores across almost all evaluation metrics i.e., Completion Rate (CR), Accuracy Once, Accuracy Range, Accuracy Periodic, and Accuracy Average.
xxModel CR (%) Accuracy (%)
xOnce Range Periodic Avg.
Llama H2O 64 45 60 27 44
SnapKV 64 50 55 26 44
MorphKV 64 50 61 24 45
Mistral H2O 71.2 57 60 32 50
SnapKV 71 55 57 36 49
MorphKVx 71.2 57 62 36 52
Qwen H2O 55 46 51 28 42
SnapKV 53 44 46 28 39
MorphKVx 51 43 68 30 47
5.2.1 Performance
Table 2 shows the performance of MorphKV over prior works. LongGenBench uses a rigorous evaluation suite to assess response quality. This includes information recall about singular instances (Accuracy Once), range of instances (Accuracy Range), periodic instances (Accuracy Periodic) and their average (Average Accuracy), while Completion Rate (CR) quantifies the percentage of tasks successfully completed. MorphKV generally outperforms or matches SnapKV, and H2O on all models and metrics. Notably, SnapKV retains all prompt tokens due to its ample cache budget. It also keeps track of every decoded token, effectively replicating Full-Attention for these tasks.
5.2.2 KV cache Sizes
Figure 8 shows the KV cache sizes for H2O, SnapKV, and MorphKV relative to Full-Attention. On average, MorphKV requires 0.55×0.$5$ while H2O and SnapKV require 1.22×1.$2$, and 5.01×5.01$\times$5.01 × the cache size of Full-Attention respectively.
Figure 8: LongGenBench: KV cache usage of H2O, SnapKV, and MorphKV relative to Full-Attention. SnapKV has an explosive KV cache size (up to 13$\times$ higher) due to the extensive number of KV pairs retained. In contrast, MorphKV maintains a constant footprint, as it retains a fixed number of recent and older tokens.
xxxxModel 2wmqa drdr hpqa mnews mfqaen mfqazh musq nqa pcnt prt qsp qms sams tqa vcs
Llama xxxSnapKV 16.0 22.0 14.9 25.6 25.4 18.7 10.7 32.2 7.6 98.4 11.7 23.1 42.9 91.7 14.2
xxxMorphKV 14.9 22.5 15.9 26.6 25.7 19.9 10.7 31.9 7.5 97.8 11.9 23.6 42.9 91.5 15.2
xxxFull-Attention 16.5 30.0 16.7 26.8 27.4 20.1 11.4 32.0 6.9 97.7 13.2 23.6 43.7 91.6 16.1
Mistral xxxSnapKV 26.6 23.7 40.5 26.0 48.8 41.3 18.3 25.6 2.5 88.6 31.0 23.8 41.9 86.3 13.5
xxxMorphKV 26.7 23.9 40.8 26.6 48.4 43.0 16.7 26.7 3.0 85.9 30.9 23.6 42.3 86.3 13.7
xxxFull-Attentionx 27.1 30.4 43.0 27.1 49.2 48.3 18.8 26.7 2.8 87.0 33.0 24.2 42.8 86.2 15.2
Phi4 xxxSnapKV 22.3 24.2 19.5 25.0 38.0 47.2 5.2 20.5 12.6 63.9 32.4 22.1 47.2 90.5 11.4
xxxMorphKV 22.6 24.1 19.3 25.5 38.2 46.4 6.2 21.0 12.6 64.3 31.2 22.4 47.6 90.6 12.3
xxxFull-Attention 22.2 29.0 19.6 25.9 38.2 48.9 6.0 20.7 11.6 63.3 33.3 22.9 48.2 90.4 13.4
Table 3: LongBench: Performance comparison of MorphKV, SnapKV, and full attention across different models. MorphKV achieves higher accuracy in most micro-benchmarks, as its KV selection algorithm minimizes redundancy and noise in the attention profile.
Figure 9: LongBench: (a) Llama3.1-8b-instruct KV cache sizes of SnapKV, and MorphKV relative to full attention. On an average, SnapKV has a KV cache size of 0.42×0.42$\times$0.42 ×, whereas MorphKV is 0.15×0.15$\times$0.15 × compared to Full-Attention (b) Average KV cache sizes of SnapKV and MorphKV relative to full attention across different models. MorphKV yields comparable performance to SnapKV at roughly 50% lower KV cache budget in a long-context setting, where the prompt is significantly larger than the response.
5.3 Long-Context: LongBench Tasks
Apart from being memory-efficient for long-response tasks, MorphKV also offers competitive performance as the state-of-the-art prompt KV compression for long-context tasks. We evaluate MorphKV, SnapKV, and Full-Attention across benchmarks in LongBench. LongBench is a comprehensive, bilingual, multitask benchmark suite used to evaluate LLMs for processing extended contexts. It comprises datasets across six task categories in English and Chinese, with an average prompt length of nearly 6K tokens. For MorphKV, we set the recent window to 32 tokens and fix its total cache capacity at 2K tokens, and use the sum() fusion. In contrast, SnapKV preserves 1024 tokens from the prefill phase and all decoded tokens across all attention heads.
5.3.1 Performance
Table 3 shows that MorphKV generally matches or outperforms SnapKV across most datasets. Notably, MorphKV consistently surpasses SnapKV for MultiNews on all models. Moreover, for tasks like Phi4-2WikiMQA, Phi4-Passage-Count, and Phi4-TriviaQA, MorphKV even exceeds Full-Attention performance while using only 20% of the memory capacity. This suggests that larger models (e.g., Phi4 with 14B parameters) can better leverage the dynamic token selection in MorphKV to capture essential information.
5.3.2 KV cache Sizes
Figure 9 compares the average KV cache memory usage of SnapKV and MorphKV relative to Full-Attention across all LongBench datasets. MorphKV achieves up to 2×2$\times$2 × memory savings over SnapKV and up to 5×5$\times$5 × over Full-Attention. Notably, for datasets like MultiNews, SnapKV requires 2×2$\times$2 × more memory than Full-Attention because it retains KV pairs across all heads, whereas MorphKV operates at just 0.4×0.4$\times$0.4 × the memory of Full-Attention, benefiting from dynamic eviction and GQA compatibility. Designed for GQA, MorphKV supports 2×2$\times$2 × more tokens while using only half the KV cache capacity of SnapKV.
Llama3.1xx 2wmqa drdr hpqa lsht mnews mfqaen mfqazh musq nqa pcnt prt qsp qms sams trec tqa vcs
max_fusedxx 14.9 21.0 14.9 33.5 25.6 24.2 17.9 9.3 32.0 8.0 97.6 10.0 23.5 42.1 46 91.8 14.5
sum_fusedxx 14.6 22.0 15.0 35.5 25.6 25.4 19.2 9.9 32.2 7.9 96.9 10.5 22.9 43.1 49 91.8 14.8
Table 4: LongBench: Llama3.1-8B-Instruct comparison of MorphKV under different fusion functions with the same cache budget.
6 Ablation studies
The design of MorphKV consists of three key design parameters, namely, fusion function used to create the attention profile, recent window size, and total KV cache budget. These hyper-parameters collectively influence the final benchmark performance with MorphKV. In this section, we analyze the impact of each hyper-parameter by experimenting with fusion functions, testing various window sizes, and substantially compressing the KV cache budget for MorphKV, while drawing comparisons with existing methods.
6.1 Impact of Fusion Function: sum() Versus max()
MorphKV considers two fusion functions for deriving the final attention profile, namely, $F_{i}$: $\textit{sum}()$ and $\textit{max}()$. In this subsection, we discuss their impact on the performance.
6.1.1 On Long-Context Tasks
We compare both $\textit{sum}()$ and $\textit{max}()$ fusion using the LongBench suite on the Llama3.1-8B model, with a recent window configuration of 32 tokens and KV cache capacity of 1K tokens. Table 4 shows per-dataset performance. On average, $\textit{max}()$ fusion outperforms $\textit{sum}()$ by about 1%, and up to 2.7% on QMSum. Datasets such as 2WikiMQA, MultiNews, Passage Count, Passage Retrieval (En), QMSum, and TriviaQA often demand sharply focused retrieval or reasoning. A single strongly attended token in these tasks can suffice to link crucial context, making $\textit{max}()$ advantageous. In contrast, $\textit{sum}()$ tends to retain past tokens which are preferred by majority of the window tokens. Consequently, $\textit{max}()$ better captures a small set of pivotal tokens spread over a large distance (longer sequence of tokens).
6.1.2 On Long-Response Tasks
xFusionx xRelevance Accuracy Coherence Clarity
Llama max 83.8 81.3 57.1 64.2
sum 89.2 81.7 63.3 71.3
Mistral max 91.7 86.7 82.9 82.1
sum 92.5 89.2 84.2 85.4
Phi4 max 62.9 79.6 68.3 72.1
sum 62.5 80 70.4 75.0
Qwen max 83.3 70.8 58.3 60.4
sum 85.4 70.4 58.3 59.1
Table 5: LongWriter: Sensitivity to $\textit{sum}()$ versus $\textit{max}()$ fusion across different models. These fusion functions dictate how attention scores are aggregated to make the final attention profile.
We similarly compare the sensitivity of $\textit{sum}()$ and $\textit{max}()$ for LongWriter tasks which contain essay-style long response prompts. For our studies, we fix the recent window to be 30 tokens, and KV cache capacity to a total of 600 tokens. As shown in Table 5, $\textit{sum}()$ fusion tends to be more effective for most models, except Qwen2.5 where $\textit{max}()$ excels in certain metrics. LongWriter tasks are typically open-ended, causing $\textit{max}()$ to emphasize specific tokens that are not always universally relevant. Conversely, $\textit{sum}()$ aggregates attention across the recent window, providing a broader (though slightly noisier) context that suits open-ended generation.
6.2 Impact of Window Size on Long-Response Tasks
We evaluate the impact of window size on the overall performance of MorphKV. Intuitively, recent-window tokens are used to select which tokens to retain from the distant past. Hence, a larger window allows capturing of more diverse information from the past.
Particularly, for LongGenBench, this effect is evident since the prompts contain lot of information which might be needed at a much later point in the generation. Hence, we run the LongGenBench suite for Llama3.1-8B, Mistral-7B and Qwen2.5 models for two variants of MorphKV, both using max fusion and a total capacity of 4K tokens. The first configuration uses a 32-token window, while the second uses a 200-token window.
xxConfig xCR(%)x Accuracy(%)
xOnce Range Periodic Avg.
Llama (32, 4K) 64 43 56 27 42
(200, 4K) 64 50 61 24 45
Mistral (32, 4K)x 71 57 60 32 50
(200, 4K)x 71 57 62 36 52
Qwen (32, 4K) 52 41 43 34 40
(200, 4K) 51 43 68 30 47
Table 6: LongGenBench: Sensitivity of evaluation metrics with window size across different models (CR: Completion Rate)
As shown in Table 6, changing the window size significantly impacts evaluation metrics for both the Llama and Mistral models. This is due to the fact that a very small window does not suffice for capturing extensive amounts of distant information present in a typical LongGenBench prompt (such as specific details about different floors in a building, specific Menu items etc.), because very small windows tend to capture local context, while failing to capture more distant context. Therefore, a larger window is more effective since it allows the model to retain diverse pieces of information even at very large response sizes. For instance, a window of size 200 lets Llama3.1 recall accurate instructions regarding the 96th floor of a building, in spite of already generating extensive descriptions of the previous 95 floors. On the other hand, with a window size of 32 tokens, the model struggles to maintain consistency with the input request, and generates generic responses after certain number of floors, thereby losing on accuracy.
Note that the Completion Rates for both window sizes are comparable. This is because the Completion Rate measures the number of times the model was able to generate what it was expected to (for example, how many floors did the model generate the floor plan for out of the requested 100 floors). This is a relatively simpler task, and a smaller window can keep track of such information (for instance, by simply retaining the floor number of the last floor it generated the plan for). Consequently, we do not observe substantial differences in the Completion Rate metric.
6.3 Robustness Against KV Cache Compression
To assess the impact of KV cache compression on MorphKV versus SnapKV, we run ablation studies on a subset of benchmarks within the LongBench suite. We record the resulting performance across Llama3.1-8B and Mistral-7B models. For both MorphKV and SnapKV, the KV cache budget is varied from 1% to 7% with respect to full attention KV cache size. MorphKV uses a window of 32 tokens, with $sum()$ as the fusion function. SnapKV also uses the same window size of 32 tokens, but maintains KV cache across all attention heads. This enables MorphKV to store 4×4$\times$4 × as many tokens as SnapKV at the same cache capacity.
Figure 10: LongBench: Average scores for (a) Llama3.1-8B and (b) Mistral-7B models, comparing MorphKV and SnapKV under different levels of KV cache compression. SnapKV degrades much more sharply with decreasing KV cache budget, while MorphKV offers stable performance even at a low KV cache budget.
Figure 10 shows the mean benchmark score (we refer to Appendex A.1 for individual benchmark scores) for both models under varying compression scenarios. At very low KV cache budgets, we observe a drop of more than $50\%$ on average between MorphKV and SnapKV, for benchmarks like NarrativeQA, this difference reaches upto $88\%$. This disparity indicates that MorphKV is significantly more effective at retaining crucial context information compared to SnapKV under tight memory constraints. Even with larger budgets, MorphKV consistently outperforms SnapKV, demonstrating the robustness and reliability of its design.
The input prompt for long-response tasks is typically very small, and SnapKV does not evict KV cache during decoding, hence we exclude a similar analysis for these tasks.
7 Conclusion
The growing memory footprint of KV caches in LLMs poses a critical bottleneck for long-context and long-response tasks. In this paper, we propose MorphKV that addresses this challenge by introducing a dynamic, correlation-aware token selection mechanism that maintains a constant-sized KV cache while preserving contextual coherence. MorphKV leverages attention profiles of recent tokens to identify and retain only the most relevant distant tokens. Our studies on long-response tasks show 52.9% memory savings and 18.2% higher accuracy on average compared to state-of-the-art prior works. Our experiments demonstrate that MorphKV scales efficiently with response length, degrading only 10% in performance even as outputs grow to 12K tokens, compared to 15–18% degradation for state-of-the-art prior works, SnapKV and H2O. Furthermore, MorphKV’s compatibility with Grouped Query Attention enables 4$\times$ greater memory efficiency than Multi Headed Attention-based approaches, making it practical for real-world deployment in chatbots, interactive assistants, and content generation systems. These advances position MorphKV as a practical inference-time solution for LLMs, balancing both accuracy and memory constraints without sacrificing the ability to capture long-range dependencies.
Acknowledgments
The authors acknowledge the Texas Advanced Computing Center (TACC) and the Center for Generative AI at the University of Texas at Austin for providing computational resources that helped develop the research results reported in this paper. Poulami Das acknowledges the generous support through the AMD endowment at the University of Texas at Austin. Prashant J. Nair is supported by Intel Transformation Server Architecture (TSA) and the Natural Sciences and Engineering Research Council of Canada (NSERC) [funding reference number RGPIN-2019-05059] Grants. We acknowledge Won Joon Yun for editorial feedback.
Appendix A Appendix
A.1 LongBench: Comparison of MorphKV and SnapKV under different KV cache budget constraints
We stress-test both MorphKV and SnapKV for robustness against KV cache compression, by varying the compression rates on LongBench. Figure 11 shows per dataset trends, we note that MorphKV consistently provides robust performance even at extreme levels of KV cache compression, while SnapKV’s performance degrades drastically at very low KV cache budget.
Figure 11: Comparison of MorphKV versus SnapKV for Llama3.1-8B, and Mistralv0.2-7B. (a)-(c) show results for Llama3.1-8B on MultifieldQA-en, Musique, and NarrativeQA respectively, while (d)-(f) corresponds to Mistralv0.2-7B evaluation for NarrativeQA, PassageRetrieval-en, and 2wikimqa respectively. MorphKV consistently beats SnapKV across varying KV cache budget sizes.
A.2 LongWriter: LLM Judge Scores for Different Models and Metrics
Table 7 provides detailed LLM Judge Scores across various criteria. All configurations use a KV cache capacity of 600 tokens, with a recent window size of 30 tokens for both SnapKV and MorphKV. MorphKV consistently outperforms H2O on Relevance, highlighting its effectiveness in retaining imporatant tokens. Qwen2.5 differs from other models as it employs GQA with 7×7$\times$7 × less heads than the default MHA configuration, potentially introducing partial information loss when MorphKV operates under fewer heads, explaining why MorphKV performs compared to SnapKV for this model.
Modelxx xRelevance Accuracy Coherence Clarity Breadth and Depth Reading Experience Total
Llama3.1-8B H2O 84.6 81.7 63.3 71.7 54.2 55.4 68.5
SnapKV 85.8 80.4 63.3 73.3 48.8 54.6 67.7
MorphKV 89.2 81.7 63.3 71.2 57.1 54.6 69.5
Full-Attention 86.2 81.7 57.9 71.2 49.2 52.5 66.5
Mistral-7B H2O 91.7 89.6 81.2 83.3 60.4 73.8 80.0
SnapKV 90.4 89.6 84.6 84.2 61.7 76.2 81.1
MorphKV 92.5 89.2 84.2 85.4 60.4 75.0 81.1
Full-Attentionxx 93.8 88.8 85.0 84.6 60.4 75.0 81.2
Phi4 H2O 59.2 77.9 63.8 70.8 40.4 57.1 61.5
SnapKV 66.7 78.3 68.3 71.2 42.5 55.8 63.8
MorphKV 62.5 80 70.42 75 41.2 58.75 64.65
Full-Attention 65.0 78.3 63.8 72.9 43.8 53.8 62.9
Qwen2.5 H2O 85.0 67.1 54.6 56.2 67.1 52.5 63.8
SnapKV 87.7 72.5 60.6 63.6 68.2 57.6 68.4
MorphKV 85.4 70.4 58.3 59.2 63.8 52.5 64.9
Full-Attention 86.4 69.5 59.8 61.0 64.4 56.4 66.2
Table 7: LongWriter: LLM Judge Scores by model across multiple metrics.